Can every vector in R4 be written as a linear combination of the columns of matrix B above? Do...

B is the coefficient matrix = 1 4 1 2

0 1 3 -4

0 2 6 7

2 9 5 -7

The part of the question that confuses me is the part that asks if the columns of B span R3 or not because I am only sure of R4.
Answer #1
Question Details

Can every vector in R4 be written as a linear combination

of the columns of matrix B above? Do the columns of B span R3?

B is the coefficient matrix = 1 4 1 2

0 1 3 -4

0 2 6 7

2 9 5 -7

The part of the question that confuses me is the part that asks if the columns of B span R3 or not because I am only sure of R4.

AT THE OUT SET PLEASE NOTE THE FOLLOWING

BASIS

TO FORM A BASIS FOR RN , WE NEED

1.WE NEED A SET OF N AND ONLY N VECTORS

2.THEY SHALL BE L.I.

TO SPAN RN ...

1.WE NEED A SET OF N OR MORE VECTORS

2.ANY N VECTORS IN THE SET SHALL BE L.I.

SO NOW WE HAVE R4..THAT IS N=4

SO WE NEED AT LEAST 4 VECTORS TO SPAN R4..WE HAVE 4 VECTORS ...OK..

NOW WE HAVE TO CHECK WHETHER THEY ARE L.I.

B=

1

4

1

2

0

1

3

-4

0

2

6

7

2

9

5

-7

NR4=R4-2R1

1

4

1

2

0

1

3

-4

0

2

6

7

0

1

3

-11

NR1=R1-4R2…NR3=R3-2R2…NR4=R4-R2

1

0

-11

18

0

1

3

-4

0

0

0

15

0

0

0

-7

WE FIND COLUMNS 1,2,3 ARE L.D.

AS THEY HAVE ZEROS IN THE III AND IV COORDINATES.

HENCE THEY CAN NOT SPAN R4…………ANSWER

THUS WE HAVE ONLY 3 INDEPENDENT COLUMN VECTORS

SAY COLUMNS.1,2 AND 4

NOW LET US CHECK WHETHER THEY SPAN R3…

WE HAVE 4 VECTORS ..

THAT IS MORE THAN 3 NEEDED TO SPAN R3….OK

3 OF THEM ARE L.I. ….OK

SO THEY SPAN R3 …ANSWER

answered by: Monique/G